# RbRL2.0: Integrated Reward and Policy Learning for Rating-based Reinforcement Learning

### v1

@@ Abstract
Reinforcement learning (RL), a common tool in decision making, learns policies from various experiences based on the associated cumulative return/rewards without treating them differently. On the contrary, humans often learn to distinguish from different levels of performance and extract the underlying trends towards improving their decision making for best performance. Motivated by this, this paper proposes a novel RL method that mimics humans’ decision making process by differentiating among collected experiences for effective policy learning. The main idea is to extract important directional information from experiences with different performance levels, named ratings, so that policies can be updated towards desired deviation from these experiences with different ratings. Specifically, we propose a new policy loss function that penalizes distribution similarities between the current policy and failed experiences with different ratings, and assign different weights to the penalty terms based on the rating classes. Meanwhile, reward learning from these rated samples can be integrated with the new policy loss towards an integrated reward and policy learning from rated samples. Optimizing the integrated reward and policy loss function will lead to the discovery of directions for policy improvement towards maximizing cumulative rewards and penalizing most from the lowest performance level while least from the highest performance level. To evaluate the effectiveness of the proposed method, we present results for experiments on a few typical environments that show improved convergence and overall performance over the existing rating-based reinforcement learning method with only reward learning.
Introduction
@@
Recent advancements in reinforcement learning (RL) (Sutton and Barto 1998) have shown promising results in solving complex robotics tasks under the assumption that proper reward functions have been designed (Tang et al. 2024). However, in many real world cases, it is often difficult and challenging to define reward functions properly. In these situations, it is often required to include humans users who either provide demonstrations in an offline setting, called learning from demonstrations (LfD), or provide feedback in an online or offline setting, called reinforcement learning from human feedback (RLHF). Popular LfD methods include inverse reinforcement learning (IRL) (Argall et al. 2009) and behavior cloning (BC) (Torabi, Warnell, and Stone 2018), which take human expert demonstrations as the input and learn policies that mimic the demonstrations without learning the reward functions. In contrast, RLHF takes short video clips or segments and asks humans to provide feedback. The typical form of feedback can take the form of preferences over segment pairs, also known as preference-based reinforcement learning (PbRL) (Christiano et al. 2017), or ratings for individual segments, also known as rating-based reinforcement learning (RbRL) (White et al. 2024). Variants of the PbRL methods include ranking-based RL and crowd-sourcing PbRL (Brown, Goo, and Niekum 2020; Chhan, Novoseller, and Lawhern 2024). RLHF methods focus on learning reward functions from the human feedback and then training policies from the learned rewards.
@@
Recent studies in the fields of LfD and RLHF have shown promising capabilities individually. However, integrating reward learning and policy learning is challenging in the preference setting since preferences are provided in the relative sense without indicating if a segment is “optimal” or “sub-optimal”. As a contrary, RbRL utilizes ratings as the feedback and hence allows the evaluation of individual samples based on their actual performance. For example, if a sample is labeled “Very Good” (or “Very Bad” respectively) which means “desired” (or “undesired” respectively) that the policy learning should mimic (or distinguish respectively). Meanwhile, the additional ratings of “Good”, “Average”, “Bad”, indicate that the policy learning should deviate it from the three classes in an descending order. In other words, the learned policy should be closest to the samples in the “Very Good” rating, while less closer to the samples in the “Good” rating, and follow a similar trend to be most different from samples in the “Very Bad” rating class. Although such a concept is intuitive in humans’ decision making process, the current RbRL approach (White et al. 2024) only used ratings to learn reward without utilizing different rating classes directly in policy learning, which is the focus of the current work.
@@
In this paper, we focus on developing a new rating-based reinforcement learning algorithm that integrates the existing reward learning and a new policy learning motivated by how humans update the decision by shaping it away from different ratings. More precisely, the differences from different ratings are inverse proportional to their rating classes. Namely, a higher rating class means a lower difference, vice versa. Towards this objective, we first propose a novel Kullback–Leibler (KL) divergence based loss function for different rating classes that penalizes the distribution similarities between the current policy and the segments in different classes. In particular, we classify trajectories not in the highest rating class as failed segments and apply the new loss function with different weights based on their rating labels. For example, when rating class is set to 4 (rating class “0”, “1”, “2” and “3”), the new loss function calculates three distinct KL divergencies between the current policy and trajectories in rating class “0”, “1”and “2” with different weights. Second, we modularize this novel loss function, making it seamlessly integrated with the existing RbRL approach (White et al. 2024) across different rating classes without changing the original settings. Third, we conduct experimental studies to evaluate the proposed approach in several environments with different complexity levels and show that our new approach can yield better performance than the existing RbRL approach.
Related Work
@@
Learning from demonstrations (LfD) has shown as an effective method to improve RL in dense-reward, sparse-reward and reward-free environments (Schaal 1996; Subramanian, Isbell Jr, and Thomaz 2016). Methods, such as DQfD (Hester et al. 2018), DDPGfD (Vecerik et al. 2017) and work in (Nair et al. 2018), have shown significant improvement in RL by employing expert demonstrations to guide the policy searching. Specifically, DQfD and DDPGfD use pre-collected expert demonstrations as replay experiences participating in the policy updating process for Deep Q Network (DQN) and Deep Deterministic Policy Gradient (DDPG) (Mnih et al. 2013; Lillicrap et al. 2015). In (Nair et al. 2018), a separate buffer is created to store pre-collect demonstrations which are then used with behavior cloning by penalizing the dissimilarity between the current agent’s behavior and the demonstrations. However, expert demonstrations are costly available in practical, which leaves a space for researchers to explore the potential of suboptimal demonstrations or failed experiences in optimizing RL. Noisy demonstrations, with both expert and suboptimal demonstrations, are used as pre-trained data in NAC to initialize a policy and then refine this policy by interacting with environmental rewards (Gao et al. 2018).
@@
Failed experiences are used as negative examples to guide the agent’s exploration direction in (Wu et al. 2024). By penalizing the similarity between the current agent’s behaviors and those from failed experiences, this method significantly improves offline RL method via overcoming the difficulty in exploration under sparse reward settings. However, failed experiences can differ from each other as they are defined as behaviors that are not optimal. Equally treating all of these experiences could lead to unstable exploration. Existing works such as PbRL (Christiano et al. 2017) and RbRL (White et al. 2024) label the on-policy segments to indicate their different levels of performance. In reward-free environments, these labels are used to infer reward functions for optimizing RL policies. However, in both PbRL and RbRL, unpreferred segments or those with lower ratings are often abundant, leading to an underutilization of the potential value in failed experiences. To address this, our approach aims to optimize the policy by effectively leveraging failed experiences at different levels.
Preliminaries and Background
Problem Formulation
@@
Figure 1: A schematic illustration of the proposed RbRL2.0 and its relationship with RbRL.
Rating-Based Reinforcement Learning
Due to the lack of rewards, the rating-based reinforcement learning (RbRL) (White et al. 2024) learns a reward model $\hat{r}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ that predicts reward $\hat{r}(s,a)$ for each state-action pair during interaction. Given a length-j segment $\sigma=(s_{1},a_{1},...,s_{j},a_{j})$, the cumulative discounted reward $\hat{R}(\sigma):=\sum^{j}_{t}=1\gamma^{t-1}\hat{r}(s_{t},a_{t})$ based on $\hat{r}$ provides an estimated cumulative reward.
@@
[…]
@@
Rating-induced Policy Loss
The new loss function derived from the ratings can be divided into two components. As described in Figure 1, the first component is the classic loss function for gradient-based RL algorithms based on the learned reward from the rating-based reinforcement learning algorithm (White et al. 2024). The second component is our proposed new loss to penalize the similarities between the samples from the current policies and the samples from different rating classes. Since different rating classes contain samples with different performance levels, the weights used in the penalty should be different for different rating classes. The main idea of the new loss component is to use the Kullback–Leibler (KL) divergence to quantify the similarities ( equivalently, differences) with an descending weight for rating classes from low to high, which ensures more difference from low rating classes, namely, low performance levels, and less difference from high rating classes, namely, high performance levels. It is worth emphasizing that samples with high performance levels are still undesired but show better performance than those with low performance levels. Hence, all samples used in the new loss are considered failure, but with different performance levels characterized by ratings.
Overall Loss Function
@@
$-\nabla_{\theta}\sum_{i=0}^{n-2}\omega_{i}D_{KL}(D_{i}\parallel D_{\pi_{\theta}}),$
where the first component is the classic loss based on the learned reward and the second component is the new loss, $\sigma_{\theta}$ includes the trajectories sampled from the current policy at each training batch, $D_{KL}$ denotes the KL divergence between two different distributions (please refer to Definition 1 below), $\omega_{i}$ is the weight applied to the KL divergence in a descending order as rating level moves from lowest to highest, and $D$ denotes the distribution. This new policy gradient loss effectively penalizes the similarities between the current policy and different rating levels of the failed experiences such that the policy is updated in the direction of continuously improving the performance thanks to the descending order for the weight $\omega_{i}$.
To facilitate the understanding of this new loss, we will now provide some details. In the context of the new loss function, we utilize all, except the highest, rating classes. For example, consider the case of 4 rating classes, namely “very bad”, “bad”, “good” and “very good”, whose ratings are “0”, “1”, “2” and “3”, respectively. The segments rated as “very bad” are stored in buffer “0”, namely, $R_{0}$. The segments rated as “bad” in buffer “1”, namely, $R_{1}$. The segments rated as “good” in buffer “2”, namely $R_{2}$. Finally, the segments rated as “very good” in buffer “3”, namely, $R_{3}$. All trajectories in $R_{0}$, $R_{1}$ and $R_{2}$ are used in (5). In other words, the rating classes used in (5) include class 00 to class $n-2$, which is reflected in the second term of (5) with the summation computed for $i=0$ to $i=2$.
In RL, a policy is initialized with a probability density function, also referred to as a distribution, mapping from the states input and the actions output. The main goal of the RL agent is to learn an optimal policy that maximizes the cumulative rewards. In the context of rating-based reinforcement learning, the cumulative rewards are based on the estimated rewards since the reward is unknown or does not exist. Besides the class loss term, the proposed new loss term penalizes the deviation between the current policy and different rating levels of failed examples. To this end, we employ the multivariate Gaussian distribution (Goodman 1963) to represent trajectories in each rating class, and use KL divergence (Hershey and Olsen 2007) to measure the distribution-wise difference between the current policy and each failure distribution. The main motivation behind this new loss term is to provide a direct policy shaping mechanism towards exploring areas that are different from various rating levels of the failed samples at different penalty weights. In other words, the RL agent will be pushed away from samples and its neighboring regions at different performance levels. The lower the performance level is, the larger the push-away force will be applied.
Policy Loss Based on KL divergence
Definition 1.
$D_{KL}(P\parallel Q)$ $=\frac{1}{2}(Tr(\Sigma^{-1}_{q}\Sigma_{p})+(\mu_{p}-\mu_{q})^{T}\Sigma^{-1}_{q}(\mu_{q}-\mu_{p})$
$-k+ln(\frac{det(\Sigma_{q})}{det(\Sigma_{p})})),$ (6)
where, $Tr(\cdot)$ is the trace of a given matrix, $det(\cdot)$ represents the determination of a given matrix, and $k$ is the number of features.
Definition 1 provides a measure of how different the two distributions $P$ and $Q$ are. For example, we have a set of failed trajectories $\sigma_{i}=((s_{i_{0}}^{0},a_{i_{0}}^{0},s_{i_{0}}^{1},a_{i_{0}}^{1},...),...,(s_{i_{m}}^{0},a_{i_{m}}^{0},s_{i_{m}}^{1},a_{i_{m}}^{1},...))$ in rating class i containing $m$ trajectories, where $i$ is not the highest rating class. Another set of trajectories $\sigma_{\pi_{\theta}}=((s_{\pi}^{0},a_{\pi}^{0},s_{\pi}^{1},a_{\pi}^{1},...),...,(s_{\pi}^{0},a_{\pi}^{0},s_{\pi}^{1},a_{\pi}^{1},...))$ sampled from the current policy $\pi$ parameterized by $\theta$ represents the behaviors of the RL agent in the current training batch. According to Definition 1, the KL divergency between the distribution of rating class i and the distribution of the current policy can be formulated as
$D_{KL}(D_{i}\parallel D_{\pi})=$ $\frac{1}{2}(Tr(\Sigma^{-1}_{D_{\pi}}\Sigma_{D_{i}})$
$+(\mu_{D_{i}}-\mu_{D_{\pi}})^{T}\Sigma^{-1}_{\pi}(\mu_{\pi}-\mu_{D_{i}})$
$+ln(\frac{det(\Sigma_{D_{\pi}})}{det(\Sigma_{D_{i}})})).$ (7)
Consider the case of 4 different rating classes (“0”, “1”, “2” and “3”). Following the computation of the KL divergence between the current policy and the samples in rating classes 0, 1, and 2 (except class 3, which is the highest rating class and hence not included as explained in the Subsection “Overall Loss Function”), the overall loss function can be written as
$\nabla_{\theta}J(\pi_{\theta})=$ $\mathbb{E}_{\pi_{\theta}}[\nabla_{\theta}\log(\pi_{\theta})\hat{R}(\sigma_{\theta})]$
$-(\omega_{0}D_{KL}(D_{0}\parallel D_{\pi_{\theta}})$
$+\omega_{1}D_{KL}(D_{1}\parallel D_{\pi_{\theta}})$
$+\omega_{2}D_{KL}(D_{2}\parallel D_{\pi_{\theta}})),$ (8)
@@
where $\hat{R}(\sigma_{\theta})$ represents the cumulative predicted rewards of the trajectories sampled by the current policy $\pi_{\theta}$, $\omega_{0}$, $\omega_{1}$ and $\omega_{2}$ represents the weights for KL divergencies between rating classes “0”, “1” and “2” and the current policy $\pi_{\theta}$, respectively. The weights are assigned in the descending order, namely $\omega_{0}>\omega_{1}>\omega_{2}$, such that current policy $\pi_{\theta}$ is pushed away from the distribution of rating class “0” more than distributions of rating classes “1” and “2”.
The pseudocode of the proposed new approach is given in Algorithm 1.
Algorithm 1 RbRL2.0
@@
rating classes $n$, rating buffers $R_{1},...,R_{n}$, initial reward predictor $\hat{R}$, KL divergency weight $\omega_{i}$, total training cycles $T$, total training cycle $M$ for $\hat{R}$
@@
$\nabla_{\theta}J(\pi_{\theta})$ $=\mathbb{E}_{\pi_{\theta}}[\nabla_{\theta}\log(\pi_{\theta})\hat{R}(\sigma_{\theta})]$
$-\nabla_{\theta}\sum_{i=0}^{n-2}\omega_{i}D_{KL}(D_{i}\parallel D_{\pi_{\theta}})$
@@
Bridge Between Machine Learning and Cognitive Science
Fig. 1 is a flowchart of the proposed integrated reward and policy learning approach. This approach bridges machine learning and cognitive science, where reinforcement learning policy learning is enabled through a human-inspired rating mechanism. In particular, the rating mechanism is used to classify samples based on their different levels of performance, which can provide AI agent to learn inherent reward functions as well as guide the search of policies that deviate from these samples differently based on their performance levels.
From the machine learning perspective, the proposed approach creates a new human cognitive-based mechanism, namely, rating, into the training of reinforcement learning algorithms. Moreover, we explore the value of rated samples by following humans’ learning from failed experience directly and indirectly. The second term in the proposed policy loss in (5) directly follows the concept of “learning to differentiate” inherent in humans’ cognitive learning from failure. The first term in the proposed policy loss in (5) indirectly impacts the policy learning by maximizing the estimated reward that can best reflect the rated samples.
From the cognitive science perspective, the proposed approach provides a mathematical approach that allows AI agents to reason its actions following a similar concept of human minds. In particular, the benefits of the proposed method shown in the “Experiments and Results” Section demonstrate the value of uncovering cognitive science for their adoption in machine learning to achieve unprecedented performance via harnessing human knowledge.
Experiments and Results

### v2

# Human-Inspired Multi-Level Reinforcement Learning

@@ Abstract
Reinforcement learning (RL), a common tool in decision making, learns control policies from various experiences based on the associated cumulative return/rewards without treating them differently. Humans, on the contrary, often learn to distinguish from discrete levels of performance and extract the underlying insights/information (beyond reward signals) towards their decision optimization. For instance, when learning to play tennis, a human player does not treat all unsuccessful attempts equally. Missing the ball completely signals a more severe mistake than hitting it out of bounds (although the cumulative rewards can be similar for both cases). Learning effectively from multi-level experiences is essential in human decision making. This motivates us to develop a novel multi-level RL method that learns from multi-level experiences via extracting multi-level information. At the low level of information extraction, we utilized the existing rating-based reinforcement learning (white2024rating) to infer inherent reward signals that illustrate the value of states or state-action pairs accordingly. At the high level of information extraction, we propose to extract important directional information from different-level experiences so that policies can be updated towards desired deviation from these different levels of experiences. Specifically, we propose a new policy loss function that penalizes distribution similarities between the current policy and different-level experiences, and assigns different weights to the penalty terms based on the performance levels. Furthermore, the integration of the two levels towards multi-level RL guides the agent toward policy improvements that benefit both reward improvement and policy improvement, hence yielding a similar learning mechanism as humans. To evaluate the effectiveness of the proposed method, we present results for experiments on a few typical environments that show its advantages over the existing rating-based reinforcement learning (white2024rating), where a single reward learning was used.
1 Introduction
@@
Recent advancements in reinforcement learning (RL) (sutton1998introduction) have shown promising results in solving complex robotics tasks under the assumption that proper reward functions have been designed (tang2024deepreinforcementlearningrobotics). However, in many real world cases, it is often difficult and challenging to define reward functions properly. In these situations, it is often required to include humans users who either provide demonstrations in an offline setting, called learning from demonstrations (LfD), or provide feedback in an online or offline setting, called reinforcement learning from human feedback (RLHF). Popular LfD methods include inverse reinforcement learning (IRL) (argall2009survey) and behavior cloning (BC) (torabi2018behavioral), which take human expert demonstrations as the input and learn policies that mimic the demonstrations without learning the reward functions. In contrast, RLHF takes short video clips or segments and asks humans to provide feedback. The typical form of feedback can take the form of preferences over segment pairs, also known as preference-based reinforcement learning (PbRL) (christiano2017deep), or ratings for individual segments, also known as rating-based reinforcement learning (RbRL) (white2024rating). Variants of the PbRL methods include ranking-based RL and crowd-sourcing PbRL (brown2020better; chhan2024crowd). RLHF methods focus on learning reward functions from the human feedback and then training policies from the learned rewards.
@@
Recent studies in the fields of LfD and RLHF have shown promising capabilities individually. However, jointly learning reward learning and policy learning is challenging in the preference setting since preferences are provided in the relative sense without indicating if a segment is “optimal" or “sub-optimal". As a contrary, RbRL utilizes ratings as the feedback and hence allows the evaluation of individual samples based on their multi-level performance. For example, if a sample is labeled “Very Good” (or “Very Bad” respectively) which means “desired” (or “undesired” respectively) that the policy learning should mimic (or distinguish respectively). Meanwhile, the additional ratings of “Good”, “Average”, “Bad”, indicate that the policy learning should deviate it from the three classes in a descending order from humans’ cognitive perspective. In other words, the learned policy should be closest to the samples in the “Very Good” rating, while less closer to the samples in the “Good” rating, and follow a similar trend to be most different from samples in the “Very Bad” rating class. Although such a concept is intuitive in humans’ decision making process, the current RbRL approach (white2024rating) only used ratings to learn reward without utilizing different levels of performance behind different rating classes in high-level policy direction learning, which is the focus of the current work.
@@
In this paper, we propose a novel multi-level reinforcement learning algorithm that leverages human ratings in two levels, namely, low-level reward learning and high-level policy direction learning that guides the agent to systematically deviate from experiences of different rating levels in a human-inspired way. More precisely, the differences from different ratings, or equivalently performance levels, are inverse proportional to their rating classes. Namely, a higher rating class means a lower difference, vice versa. Towards this objective, we first propose a novel Kullback–Leibler (KL) divergence based loss function for different rating classes that penalizes the distribution similarities between the current policy and the segments in different classes. In particular, we classify trajectories not in the highest rating class as failed segments and apply the new loss function with different weights based on their rating labels. For example, when rating class is set to 4 (rating class “0”, “1”, “2” and “3”), the new loss function calculates three distinct KL divergencies between the current policy and trajectories in rating class “0”, “1”and “2” with different weights. Second, we design this loss function in a modular and flexible manner, allowing it to operate across all rating classes while preserving the original RbRL framework (white2024rating), so that the multi-level information derived from the rated data can be applied seamlessly without altering existing training procedures. Third, we conduct experimental studies to evaluate the proposed approach in several environments with different complexity levels and show that the proposed algorithm can yield better performance than the existing RbRL approach.
2 Related Work
@@
Learning from demonstrations (LfD) has shown as an effective method to improve RL in dense-reward, sparse-reward and reward-free environments (schaal1996learning; subramanian2016exploration). Methods, such as DQfD (hester2018deep), DDPGfD (vecerik2017leveraging) and work in (nair2018overcoming), have shown significant improvement in RL by employing expert demonstrations to guide the policy searching. Specifically, DQfD and DDPGfD use pre-collected expert demonstrations as replay experiences participating in the policy updating process for Deep Q Network (DQN) and Deep Deterministic Policy Gradient (DDPG) (mnih2013playing; lillicrap2015continuous). In (nair2018overcoming), a separate buffer is created to store pre-collect demonstrations which are then used with behavior cloning by penalizing the dissimilarity between the current agent’s behavior and the demonstrations. However, expert demonstrations are costly available in practical, which leaves a space for researchers to explore the potential of suboptimal demonstrations or failed experiences in optimizing RL. Noisy demonstrations, with both expert and suboptimal demonstrations, are used as pre-trained data in NAC to initialize a policy and then refine this policy by interacting with environmental rewards (gao2018reinforcement).
@@
Failed experiences are used as negative examples to guide the agent’s exploration direction in (wu2024offline). By penalizing the similarity between the current agent’s behaviors and those from failed experiences, this method significantly improves offline RL method via overcoming the difficulty in exploration under sparse reward settings. However, failed experiences can differ from each other as they are defined as behaviors that are not optimal wu2025robust. Equally treating all of these experiences could lead to unstable exploration. Existing works such as PbRL (christiano2017deep) and RbRL (white2024rating) label the on-policy segments to indicate their different levels of performance. In reward-free environments, these labels are used to infer reward functions for optimizing RL policies. However, in both PbRL and RbRL, unpreferred segments or those with lower ratings are often abundant, leading to an underutilization of the potential value in failed experiences. To address this, our approach aims to optimize the policy by effectively leveraging failed experiences at different levels.
3 Preliminaries and Background
3.1 Problem Formulation
@@
Figure 1: A schematic illustration of the proposed algorithm and its relationship with RbRL.
3.2 Rating-Based Reinforcement Learning
Due to the lack of rewards, the rating-based reinforcement learning (RbRL) (white2024rating) learns a reward model $\hat{r}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ that predicts reward $\hat{r}(s,a)$ for each state-action pair during interaction. Given a length-j segment $\sigma=(s_{1},a_{1},...,s_{j},a_{j})$, the cumulative discounted reward $\hat{R}(\sigma):=\sum^{j}_{t}=1\gamma^{t-1}\hat{r}(s_{t},a_{t})$ based on $\hat{r}$ provides an estimated cumulative reward.
@@ 3.2 Rating-Based Reinforcement Learning
where $\bar{R}_{i}$ and $\bar{R}_{i+1}$ are the lower and upper bound value for the $i\textsuperscript{th}$ class respectively. The reward predictor $\hat{r}$ was trained by minimizing the cross-entropy loss given by
@@
4 Multi-Level Reinforcement Learning
Figure 1 shows the schematic structure of the new multi-level reinforcement learning approach. Its loss function can be divided into two components. The first component is the classic loss function for gradient-based RL algorithms based on the learned reward from the rating-based reinforcement learning algorithm (white2024rating). The second component is a new loss that penalizes the similarities between the samples from the current policies and the samples from different rating classes. Since different rating classes contain samples with different performance levels, the weights used in the penalty should be different for different rating classes. The main idea of the new loss component is to use the Kullback–Leibler (KL) divergence to quantify the similarities (equivalently, differences) with an descending weight for rating classes from low to high, which ensures more difference from low rating classes, namely, low performance levels, and less difference from high rating classes, namely, high performance levels. It is worth emphasizing that samples with high performance levels are still undesired but show better performance than those with low performance levels. Hence, all samples used in the new loss are considered failure, but with different performance levels characterized by different rating classes.
4.1 Overall Loss Function
@@
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\pi_{\theta}}[\nabla_{\theta}\log(\pi_{\theta})\hat{R}(\sigma_{\theta})]-\nabla_{\theta}\sum_{i=0}^{n-2}\omega_{i}D_{KL}(D_{i}\parallel D_{\pi_{\theta}}),$ (5)
where the first component is the classic loss based on the learned reward and the second component is the new loss, $\sigma_{\theta}$ includes the trajectories sampled from the current policy at each training batch, $D_{KL}$ denotes the KL divergence between two different distributions (please refer to Definition 4.1 below), $\omega_{i}$ is the weight applied to the KL divergence in a descending order as rating level moves from lowest to highest, while $D_{i}$ and $D_{\pi_{\theta}}$ represent the trajectory distributions of rating class $i$ and the current policy, respectively. This new policy gradient loss effectively penalizes the similarities between the current policy and different rating levels of the failed experiences such that the policy is updated in the direction of continuously improving the performance thanks to the descending order for the weight $\omega_{i}$.
To facilitate the understanding of this new loss, we will now provide some details. In the context of the new loss function, we utilize all, except the highest, rating classes. For example, consider the case of 4 rating classes, namely “very bad", “bad", “good" and “very good", whose ratings are “0", “1", “2" and “3", respectively. The segments rated as “very bad" are stored in buffer “0", namely, $R_{0}$. The segments rated as “bad" in buffer “1", namely, $R_{1}$. The segments rated as “good" in buffer “2", namely $R_{2}$. Finally, the segments rated as “very good" in buffer “3", namely, $R_{3}$. All trajectories in $R_{0}$, $R_{1}$ and $R_{2}$ are used in (5). In other words, the rating classes used in (5) include class 0 to class n2n-2, which is reflected in the second term of (5) with the summation computed for $i=0$ to $i=2$.
In RL, a policy is initialized with a probability density function, also referred to as a distribution, mapping from the states input and the actions output. The main goal of the RL agent is to learn an optimal policy that maximizes the cumulative rewards. In the context of rating-based reinforcement learning, the cumulative rewards are based on the estimated rewards since the reward is unknown or does not exist. Besides the class loss term, the proposed new loss term penalizes the deviation between the current policy and different rating levels of failed examples. To this end, we employ the multivariate Gaussian distribution (goodman1963statistical) to represent trajectories in each rating class, and use KL divergence (hershey2007approximating) to measure the distribution-wise difference between the current policy and each failure distribution. The main motivation behind this new loss term is to provide a direct policy shaping mechanism towards exploring areas that are different from various rating levels of the failed samples at different penalty weights. In other words, the RL agent will be pushed away from samples and its neighboring regions at different performance levels. The lower the performance level is, the larger the push-away force will be applied.
4.2 Policy Loss Based on KL divergence
Definition 4.1.
$D_{KL}(P\parallel Q)=\frac{1}{2}(\mathrm{Tr}(\Sigma^{-1}_{q}\Sigma_{p})+(\mu_{p}-\mu_{q})^{T}\Sigma^{-1}_{q}(\mu_{q}-\mu_{p})-k+\ln(\frac{\det(\Sigma_{q})}{\det(\Sigma_{p})})),$ (6)
where, $\mathrm{Tr}(\cdot)$ is the trace of a given matrix, $\det(\cdot)$ represents the determination of a given matrix, and $k$ is the number of features.
Definition 4.1 provides a measure of how different the two distributions $P$ and $Q$ are. For example, we have a set of failed trajectories $)\sigma_{i}=((s_{i_{0}}^{0},a_{i_{0}}^{0},s_{i_{0}}^{1},a_{i_{0}}^{1},...),...,(s_{i_{m}}^{0},a_{i_{m}}^{0},s_{i_{m}}^{1},a_{i_{m}}^{1},$...)) in rating class i containing $m$ trajectories, where $i$ is not the highest rating class. Another set of trajectories $)\sigma_{\pi_{\theta}}=((s_{\pi}^{0},a_{\pi}^{0},s_{\pi}^{1},a_{\pi}^{1},...),...,(s_{\pi}^{0},a_{\pi}^{0},s_{\pi}^{1},a_{\pi}^{1},$...)) sampled from the current policy $\pi$ parameterized by $\theta$ represents the behaviors of the RL agent in the current training batch. According to Definition 4.1, the KL divergency between the distribution of rating class i and the distribution of the current policy can be formulated as
$D_{KL}(D_{i}\parallel D_{\pi})=\frac{1}{2}(\mathrm{Tr}(\Sigma^{-1}_{D_{\pi}}\Sigma_{D_{i}})+(\mu_{D_{i}}-\mu_{D_{\pi}})^{T}\Sigma^{-1}_{\pi}(\mu_{\pi}-\mu_{D_{i}})+\ln(\frac{\det(\Sigma_{D_{\pi}})}{\det(\Sigma_{D_{i}})})).$ (7)
Consider the case of 4 different rating classes (“0", “1", “2" and “3"). Following the computation of the KL divergence between the current policy and the samples in rating classes 0, 1, and 2 (except class 3, which is the highest rating class and hence not included as explained in the Subsection “Overall Loss Function”), the overall loss function can be written as
@@ 4.2 Policy Loss Based on KL divergence
$-\left[\omega_{0}D_{KL}(D_{0}\parallel D_{\pi_{\theta}})+\omega_{1}D_{KL}(D_{1}\parallel D_{\pi_{\theta}})+\omega_{2}D_{KL}(D_{2}\parallel D_{\pi_{\theta}})\right],$
@@
where $\hat{R}(\sigma_{\theta})$ represents the cumulative predicted rewards of the trajectories sampled by the current policy $\pi_{\theta}$, $\omega_{0}$, $\omega_{1}$ and $\omega_{2}$ represents the weights for KL divergencies between rating classes “0”, “1” and “2” and the current policy $\pi_{\theta}$, respectively. The weights are assigned in the descending order, namely $\omega_{0}>\omega_{1}>\omega_{2}$, such that current policy $\pi_{\theta}$ is pushed away from the distribution of rating class “0" more than distributions of rating classes “1” and “2”. The pseudocode of the proposed new approach is given in Algorithm 1.
Algorithm 1
@@
Input: rating classes $n$, rating buffers $R_{1},...,R_{n}$, initial reward predictor $\hat{r}$, KL divergency weight $\omega_{i}$, total training cycles $T$, total training cycle $M$ for $\hat{r}$
@@
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\pi_{\theta}}[\nabla_{\theta}\log(\pi_{\theta})\hat{R}(\sigma_{\theta})]-\nabla_{\theta}\sum_{i=0}^{n-2}\omega_{i}D_{KL}(D_{i}\parallel D_{\pi_{\theta}})$
@@
5 Experiments and Results
To evaluate the effectiveness of our proposed method, we compare the new method, labeled as RbRL-KL, with RbRL across 6 DeepMind Control environments (tassa2018deepmind), namely, Cartpole-balance, Ball-in-cup, Finger-spin, HalfCheetah, Walker and Quadruped. These environments are characterized with continuous state and action spaces, and each vary in complexity. Specifically, Cartpole-balance is characterized by a simple 5-dimensional state space for cart and pole dynamics, and a 1-dimensional action space, representing discrete control forces. Ball-in-cup has an 8-dimensional state space capturing the relative positions and velocities of the involved objects, and a 2-dimensional action space controlling the cup’s motion. Finger-spin is characterized with a 12-dimensional state space representing the finger and object dynamics, and a 2-dimensional action space controlling the finger’s movement. HalfCheetah has a 17-dimensional state space, capturing joint positions, velocities, and body orientation, and a 6-dimensional action space, which corresponds to the torques applied to each joint. Walker has a 24-dimensional state space representing joint angles, velocities, and torso orientation, and a 6-dimensional action space, controlling forces applied to each limb. Quadruped is a more complex environment which characterized with a 78-dimensional state space, including joint positions, velocities, and full body orientation, along with a 12-dimensional action space, enabling torque-based control over each leg joint. Among these environments, Cartpole-balance is the simplest environment while Quadruped is the most complex one. We focus on evaluating and understanding how RbRL-KL performs across different levels of complexity. While these environments include built-in reward functions, we deliberately avoid using them to preserve a reward-free testing setup. Instead, the original reward functions are used to purely evaluate the performance of the trained policies.
(a) Cartpole-balance
(b) Ball-in-cup
(c) Finger-spin
(d) HalfCheetah
(e) Walker
(f) Quadruped
Figure 2: Learning curves of different algorithms across six environments. The plots show the mean (solid line) with the standard error (shaded area) over 10 runs.
Table 1: Hyperparameters used in experiments.